#1.) In a genetics experiment on​ peas, one sample of offspring contained 419 green peas and 11 yellow peas. Based on those​ results, estimate the probability of getting an offspring pea that is green. Is the result reasonably close to the value of 3/4 that was​ expected?

#2.) A modified roulette wheel has 44 slots. One slot is​ 0, another is​ 00, and the others are numbered 1 through 42​, respectively. You are placing a bet that the outcome is an even number.​ (In roulette, 0 and 00 are neither odd nor​ even.)
Answer #1

(1)

Green = 419

Yellow = 11

Total = 430

Estimate of the probability of getting Green = 419/430 = 0.9744

So,

Answer is:

0.9744

The result = 0.9744 is NOT  reasonably close tothe value of 3/4 = 0.75 that wasexpected.

(2)

Total number of slots = 44

Number of even number = 21

Probability of winning = P(even) = 21/44 = 0.4773

So,

Answer is:

0.4773